OpenStudy (anonymous):

graph the system of linear inequalities Y>x+4 x<_1

OpenStudy (anonymous):

 Ok there is the graph of the two inequalities. The shading sucks. The point is that y>x+4 is y values above and not including the line y=x+4 and \[x\le 1\] is a vertical line at x=4 and everything to the left. Both would be the parts that both shadings hit.

OpenStudy (anonymous):

what are the points u graphed

OpenStudy (anonymous):

For y>x+4 you could put points (0,4) and (-4,0) and draw/connect those lines. For x less or equal to 1 (1,0) and (1,5) would work and draw and connect those. Shade above y>x+4 since this means the y values above the graph Shade to the left of x less or equal to 1 line since these are the x values less than 1 Shade darkly the two shadings that overlap and thats your answer. Remember that y>x+4 does not include the line while x<= 1 does include the line

OpenStudy (anonymous):

thanks so much may u help me with another problem?

OpenStudy (anonymous):

Sure:)

OpenStudy (anonymous):

ok find the minimum and maximum valuessof the objective function subject to the givin constraints Objective function:c=x+y Constraints: X>_ 0 y>_0 y<_ 1/2x+2

OpenStudy (anonymous):

Ok you need to repeat what we did before with \[x\ge 0\] Is a vertical line at x=0 and you would shade any x value greater or equal than 0 \[y\ge 0\] Is a horizontal line and you would shade any y value greater or equal than 0 and \[y\le 0.5x+2\] is a line between points (0,2) and (-4,0) and you would shade any y value LESS than or equal to the line

OpenStudy (anonymous):

well this one does not need to be graphed but how do u solve it?

OpenStudy (anonymous):

This is not my strength, but I would think that you could do a substitution. You know that y<=0.5x+2 so C<= x+0.5x+2 c<=1.5x+2 when x=0 c=2, so that would be a minimum. I don't think there is a max though